# DeepTV: A neural network approach for total variation minimization

### v1

@@ 3.2 $\Gamma$-convergence
$\mathcal{E}_{n}(u):=\begin{cases}\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u)&\text{if }\ u\in\mathcal{H}_{c_{n}},\\ +\infty&\text{otherwise},\end{cases}$ (12)
@@
Then there exists a sequence $(c_{n})_{n}\subset\mathbb{R}$ such that the sequence $(\mathcal{E}_{n})_{n}$ $\Gamma$-converges to $\mathcal{E}$ (shortly $\operatorname*{\Gamma\mbox{-}\lim}_{n\to\infty}\mathcal{E}_{n}=\mathcal{E}$) with respect to the $L^{1}(\Omega)$ topology.
@@
Note that for any $c\geq 0$ we have $\mathcal{H}_{c}\subseteq W^{1,2}(\Omega)$. We check the two conditions in definition 3.6:
@@
    Let $(\mathcal{T}_{h_{n}})_{n}$ be a sequence of nested regular meshes of simplices with mesh size $h_{n}$; an example for $d=2$ (two dimensions) is shown in figure 1. Then for all $v\in C^{\infty}(\Omega)$, we can construct a continuous piecewise affine function $v_{h_{n}}$, whereby $v_{h_{n}}$ coincides with $v$ in every node of $\mathcal{T}_{h_{n}}$. Note that any continuous piecewise affine function can be represented by a ReLU-NN [7, Theorem 2.1]. Hence for any $n\in\mathbb{N}$ there exists a $c_{n}>0$ such that $v_{h_{n}}\in\mathcal{H}_{c_{n}}$. We choose $c_{n}$ such that $c_{n}\geq c_{n-1}$ which yields the nested property $\mathcal{H}_{c_{n-1}}\subseteq\mathcal{H}_{c_{n}}$. We define the operator $\Pi_{h_{n}}:C^{\infty}(\Omega)\to\mathcal{H}_{c_{n}}$ with $v\mapsto v_{h_{n}}=\Pi_{h_{n}}v$. By [17, Theorem 3.1.5] or [23, Corollary 1.109] there are $C_{1},C_{2}>0$ such that
@@
    $\Pi_{h_{\tilde{n}_{k}}}v_{n_{k}}\in\mathcal{H}_{c_{\tilde{n}_{k}}}$
@@
    Without loss of generality let $\tilde{n}_{k+1}>\tilde{n}_{k}$. Then by the nested property of $\mathcal{H}_{c_{n}}$ we obtain
    $\Pi_{h_{\tilde{n}_{k}}}v_{n_{k}}\in\mathcal{H}_{c_{n}}\quad\text{for all }n\geq\tilde{n}_{k}.$
    We construct a sequence $u_{n}\in\mathcal{H}_{c_{n}}$ such that
@@
    Then we have that $u_{n}\to u$ in $L^{1}(\Omega)$ and $\lim_{n\to\infty}\operatorname{TV}(u_{n})\to\operatorname{TV}(u)$, where we used that $\int_{\Omega}|\nabla v|_{2}\text{d}x=\operatorname{TV}(v)$ for all $v\in W^{1,1}(\Omega)$ [8, Section 10.1] and $\mathcal{H}_{c}\subset W^{1,1}(\Omega)$ for all $c\geq 0$. This together with the continuity of $T$ and $u\mapsto\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}$ yields
@@
Since $\mathcal{E}$ is coercive on $BV(\Omega)$, there exists a constant $a>0$ such that $S:=\{u\in BV(\Omega)\colon\mathcal{E}(u)\leq a\}$ is bounded in $BV(\Omega)$. In particular we choose $a:=\mathcal{E}(0)=\mathcal{E}_{n}(0)$, $n\in\mathbb{N}$, which ensures that any minimizer of $(\mathcal{E}_{n})_{n}$ and $\mathcal{E}$ is in $S$. In fact, for all $n\in\mathbb{N}$ a minimizer of $\mathcal{E}_{n}$ is ensured by theorem 3.4 and since $0\in\mathcal{H}_{c}$ for all $c\geq 0$ any minimizer of $\mathcal{E}_{n}$ is in $S$. Further, since $0\in BV(\Omega)$ and the $BV$-coercivity of $\mathcal{E}$ implies the existence of a minimizer of $\mathcal{E}$ in $BV(\Omega)$ [35, Theorem 3.3], also the minimizers of $\mathcal{E}$ lie in $S$.
@@
Theorem 3.7 states that there exists a sequence $(c_{n})_{n}$, but it does not provide a concrete description how to construct it. From the proof, we deduce that the sequence $(c_{n})_{n}$ must be such that $\mathcal{H}_{c_{n}}$ is sufficiently large, ensuring that the corresponding continuous piecewise affine approximation is in $\mathcal{H}_{c_{n}}$. Additionally we choose $(c_{n})_{n}$ monotonically increasing, i.e. $c_{n}\leq c_{n+1}$ for all $n\in\mathbb{N}$ which yields by the definition of $\mathcal{H}_{c}$ that $\mathcal{H}_{c_{n}}\subseteq\mathcal{H}_{c_{n+1}}$ (nested property). This allows us to construct a particular sequence $(u_{n})_{n}$. However, alternatively we could have also set $u_{k}:=\Pi_{h_{\tilde{n}_{k}}}v_{n_{k}}$ for $k\in\mathbb{N}$ and choose this as the respective converging sequence. As a consequence the nested property of $\mathcal{H}_{c}$ is not needed and hence the requirement that $(c_{n})_{n}$ is monotonically increasing can be dropped.
theorem 3.7 shows a connection between (9) and (1). In particular, for $c\to\infty$ we have that minimizers of (9) approach a minimizer of (1).
@@ 4.1 Numerical integration
To realize a numerical implementation of (9) and (11) we need to approximate its integrals. Note that, since the functions in $\mathcal{H}$ are continuous, $\operatorname{TV}(u_{\theta})=\int_{\Omega}|\nabla u_{\theta}(x)|_{2}\text{d}x$ for any $u_{\theta}\in\mathcal{H}$. Utilizing an appropriate quadrature (9) can be approximated as
$\begin{split}\min_{u_{\theta}\in\mathcal{H}_{c}}\alpha_{1}\sum_{x\in\Omega^{h}}w_{1}^{x}|(Tu_{\theta})(x)-g(x)|&+\alpha_{2}\sum_{x\in\Omega^{h}}w_{2}^{x}|(Tu_{\theta})(x)-g(x)|^{2}\\ &+\lambda\sum_{x\in\Omega^{h}}w_{\operatorname{TV}}^{x}|(\nabla u_{\theta})(x)|_{2},\end{split}$ (18)
where $\Omega^{h}\subset\Omega$ is a finite set of points and $w_{q}^{x}$, $q\in\{1,2,\operatorname{TV}\}$ are suitable quadrature weights. The superscript $x$ in $w_{q}^{x}$ emphasize that the weights might depend on $x$. Further we define $N:=|\Omega^{h}|$ the size of $\Omega^{h}$, or in other words, the number of points in $\Omega^{h}$.
@@ 4.2 Numerical realization of the total variation
$\begin{split}\min_{u_{\theta}\in\mathcal{H}_{c}}\Big{\{}E_{\theta}(u_{\theta}):=\alpha_{1}\sum_{x\in\Omega^{h}}w_{1}^{x}|(T^{h}u_{\theta}(x)-g(x)|&+\alpha_{2}\sum_{x\in\Omega^{h}}w_{2}^{x}|(T^{h}u_{\theta}(x)-g(x)|^{2}\\ &+\lambda\sum_{x\in\Omega^{h}}w_{\operatorname{TV}}^{x}|\nabla^{h}u_{\theta}(x)|_{\blacktriangleright}\Big{\}},\end{split}$ (19)
@@ 4.3 $\Gamma$-convergence
$\mathcal{E}_{n}(u):=\begin{cases}\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}+\lambda\int_{\Omega}|\nabla u(x)|\text{d}x&\text{if }u\in\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}},\\ +\infty&\text{otherwise},\end{cases}$ (24)
where $(c_{n})_{n}\subset\mathbb{R}$. Recall that for ever $c_{n}\geq 0$ the space $\mathcal{H}_{c_{n}}$ is a set of continuous piecewise affine functions. If $c_{n}$ is large enough (possibly infinite) then $\mathcal{A}_{h_{n}}\subseteq\mathcal{H}_{c_{n}}$ by [7, Theorem 2.1]. Moreover, $0\in\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}$ and hence $\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}\neq\emptyset$ for any $n\in\mathbb{N}$.
We say that the discrete functional in (19) $\Gamma$-converges to $\mathcal{E}$ defined in (13), if $(\mathcal{E}_{n})_{n}$, whereby $\mathcal{E}_{n}$ is defined in (24), $\Gamma$-converges to $\mathcal{E}$, cf. [13, Definition 4.2].
@@
Let $\mathcal{E}_{n},\mathcal{E}:L^{1}(\Omega)\to\overline{\mathbb{R}}$ be defined as in (24) and (13), respectively, and $u_{n}$ be a minimizer of $\mathcal{E}_{n}$, $n\in\mathbb{N}$. Then there exists a sequence $(c_{n})_{n}\subset\mathbb{R}$ such that $(\mathcal{E}_{n})_{n}$ $\Gamma$-converges to $\mathcal{E}$ with respect to the $L^{1}(\Omega)$ topology. Moreover, if $\mathcal{E}$ is coercive on $BV(\Omega)$, then we have that
@@
Since $\mathcal{E}$ is coercive on $BV(\Omega)$ it possesses a minimizer in $BV(\Omega)$ [35, Theorem 3.3] and the set $S:=\{u\in BV(\Omega)\colon\mathcal{E}(u)\leq a\}$ is bounded in $BV(\Omega)$, where $a:=\mathcal{E}(0)$. Since $0\in BV(\Omega)$ and $0\in\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}$ for all $n\in\mathbb{N}$, any minimizer of $(\mathcal{E}_{n})_{n}$ and $\mathcal{E}$ is in $S$. Thereby theorem 3.4 ensures a minimizer of $\mathcal{E}_{n}$. Since $\overline{S}$ is compact in $L^{1}(\Omega)$ [2, Theorem 2.5] and such that $\inf_{L^{1}(\Omega)}\mathcal{E}_{n}=\inf_{\overline{S}}\mathcal{E}_{n}$, the $\Gamma$-convergence of $(\mathcal{E}_{n})_{n}$ to $\mathcal{E}$ and [13, Theorem 1.21] yield
@@
Contrary to the proof of theorem 3.7 and [34], in the proof of theorem 4.1 we do not use the nested property of the mesh. However, this can be easily achieved by requesting that $c_{n}\geq c_{n-1}$ for all $n>1$, as in the proof of theorem 3.7.
@@
Note that if we set $c_{n}$ in (24) such that $\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}=\mathcal{A}_{h_{n}}$, then theorem 4.1 shows the $\Gamma$-convergence of a finite difference approximation of $\mathcal{E}$ to $\mathcal{E}$, as in [34]. We emphasize that showing the $\Gamma$-convergence of a finite difference approximation of $\mathcal{E}$ to $\mathcal{E}$ also seems to require a continuous approximation of a $C^{\infty}$-function. In fact, for piecewise constant approximations a $\Gamma$-convergence result cannot be shown for total variation minimization, as a counterexample presented in [10] illustrates.

### v2

@@ 3.2 $\Gamma$-convergence
$\mathcal{E}_{n}(u):=\begin{cases}\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u)&\text{if }\ u\in\mathcal{H}_{c_{n}}^{M_{n}},\\ +\infty&\text{otherwise},\end{cases}$ (12)
@@
Then there exist sequences $(c_{n})_{n}\subset\mathbb{R}$ and $(M_{n})_{n}\subset\mathbb{N}$ such that the sequence $(\mathcal{E}_{n})_{n}$ $\Gamma$-converges to $\mathcal{E}$ (shortly $\operatorname*{\Gamma\mbox{-}\lim}_{n\to\infty}\mathcal{E}_{n}=\mathcal{E}$) with respect to the $L^{1}(\Omega)$ topology.
@@
Note that for any $c\geq 0$ and $M\in\mathbb{N}$ we have $\mathcal{H}_{c}^{M}\subseteq W^{1,2}(\Omega)$. We check the two conditions in definition 3.6:
@@
    Let $(\mathcal{T}_{h_{n}})_{n}$ be a sequence of nested regular meshes of simplices with mesh size $h_{n}$; an example for $d=2$ (two dimensions) is shown in figure 1. Then for all $v\in C^{\infty}(\Omega)$, we can construct a continuous piecewise affine function $v_{h_{n}}$, whereby $v_{h_{n}}$ coincides with $v$ in every node of $\mathcal{T}_{h_{n}}$. Note that any continuous piecewise affine function can be represented by a ReLU-NN [7, Theorem 2.1]. Hence for any $n\in\mathbb{N}$ there exists a $c_{n}>0$ and an $M_{n}\in\mathbb{N}$ such that $v_{h_{n}}\in\mathcal{H}_{c_{n}}^{M_{n}}$. We choose $c_{n}$ and $M_{n}$ such that $c_{n}\geq c_{n-1}$ and $M_{n}\geq M_{n-1}$ by possibly adding new neurons to the existing hidden layers (i.e. not changing the depth of the neural networks) which yields the nested property $\mathcal{H}_{c_{n-1}}^{M_{n-1}}\subseteq\mathcal{H}_{c_{n}}^{M_{n}}$. We define the operator $\Pi_{h_{n}}:C^{\infty}(\Omega)\to\mathcal{H}_{c_{n}}^{M_{n}}$ with $v\mapsto v_{h_{n}}=\Pi_{h_{n}}v$. By [17, Theorem 3.1.5] or [23, Corollary 1.109] there are $C_{1},C_{2}>0$ such that
@@
    $\Pi_{h_{\tilde{n}_{k}}}v_{n_{k}}\in\mathcal{H}_{c_{\tilde{n}_{k}}}^{M_{\tilde{n}_{k}}}$
@@
    Without loss of generality let $\tilde{n}_{k+1}>\tilde{n}_{k}$. Then by the nested property of $\mathcal{H}_{c_{n}}^{M_{n}}$ we obtain
    $\Pi_{h_{\tilde{n}_{k}}}v_{n_{k}}\in\mathcal{H}_{c_{n}}^{M_{n}}\quad\text{for all }n\geq\tilde{n}_{k}.$
    We construct a sequence $u_{n}\in\mathcal{H}_{c_{n}}^{M_{n}}$ such that
@@
    Then we have that $u_{n}\to u$ in $L^{1}(\Omega)$ and $\lim_{n\to\infty}\operatorname{TV}(u_{n})\to\operatorname{TV}(u)$, where we used that $\int_{\Omega}|\nabla v|_{2}\text{d}x=\operatorname{TV}(v)$ for all $v\in W^{1,1}(\Omega)$ [8, Section 10.1] and $\mathcal{H}_{c}^{M}\subset W^{1,1}(\Omega)$ for all $c\geq 0$ and $M\in\mathbb{N}$. This together with the continuity of $T$ and $u\mapsto\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}$ yields
@@
Since $\mathcal{E}$ is coercive on $BV(\Omega)$, there exists a constant $a>0$ such that $S:=\{u\in BV(\Omega)\colon\mathcal{E}(u)\leq a\}$ is bounded in $BV(\Omega)$. In particular we choose $a:=\mathcal{E}(0)=\mathcal{E}_{n}(0)$, $n\in\mathbb{N}$, which ensures that any minimizer of $(\mathcal{E}_{n})_{n}$ and $\mathcal{E}$ is in $S$. In fact, for all $n\in\mathbb{N}$ a minimizer of $\mathcal{E}_{n}$ is ensured by theorem 3.4 and since $0\in\mathcal{H}_{c}^{M}$ for all $c\geq 0$ and $M\in\mathbb{N}$ any minimizer of $\mathcal{E}_{n}$ is in $S$. Further, since $0\in BV(\Omega)$ and the $BV$-coercivity of $\mathcal{E}$ implies the existence of a minimizer of $\mathcal{E}$ in $BV(\Omega)$ [35, Theorem 3.3], also the minimizers of $\mathcal{E}$ lie in $S$.
@@
Theorem 3.7 states that there exist sequences $(c_{n})_{n}$ and $(M_{n})_{n}$, but it does not provide a concrete description how to construct them. From the proof, we deduce that the sequences $(c_{n})_{n}$ and $(M_{n})_{n}$ must be such that $\mathcal{H}_{c_{n}}^{M_{n}}$ is sufficiently large, ensuring that the corresponding continuous piecewise affine approximation is in $\mathcal{H}_{c_{n}}^{M_{n}}$. Additionally we choose $(c_{n})_{n}$ and $(M_{n})_{n}$ monotonically increasing, i.e. $c_{n}\leq c_{n+1}$ and $M_{n}\leq M_{n+1}$ for all $n\in\mathbb{N}$ which yields by the definition of $\mathcal{H}_{c}^{M}$ that $\mathcal{H}_{c_{n}}^{M_{n}}\subseteq\mathcal{H}_{c_{n+1}}^{M_{n+1}}$ (nested property). This allows us to construct a particular sequence $(u_{n})_{n}$. However, alternatively we could have also set $u_{k}:=\Pi_{h_{\tilde{n}_{k}}}v_{n_{k}}$ for $k\in\mathbb{N}$ and choose this as the respective converging sequence. As a consequence the nested property of $\mathcal{H}_{c}^{M}$ is not needed and hence the requirement that $(c_{n})_{n}$ and $(M_{n})_{n}$ is monotonically increasing can be dropped.
theorem 3.7 shows a connection between (9) and (1). In particular, for $c\to\infty$ and $M\to\infty$ we have that minimizers of (9) approach a minimizer of (1).
@@ 4.1 Numerical integration
To realize a numerical implementation of (9) and (11) we need to approximate its integrals. Note that, since the functions in $\mathcal{H}^{M}$ are continuous, $\operatorname{TV}(u_{\theta})=\int_{\Omega}|\nabla u_{\theta}(x)|_{2}\text{d}x$ for any $u_{\theta}\in\mathcal{H}^{M}$. Utilizing an appropriate quadrature (9) can be approximated as
$\begin{split}\min_{u_{\theta}\in\mathcal{H}_{c}^{M}}\alpha_{1}\sum_{x\in\Omega^{h}}w_{1}^{x}|(Tu_{\theta})(x)-g(x)|&+\alpha_{2}\sum_{x\in\Omega^{h}}w_{2}^{x}|(Tu_{\theta})(x)-g(x)|^{2}\\ &+\lambda\sum_{x\in\Omega^{h}}w_{\operatorname{TV}}^{x}|(\nabla u_{\theta})(x)|_{2},\end{split}$ (18)
where $\Omega^{h}\subset\Omega$ is a finite set of points and $w_{q}^{x}$, $q\in\{1,2,\operatorname{TV}\}$, are suitable quadrature weights. The superscript $x$ in $w_{q}^{x}$ emphasize that the weights might depend on $x$. Further we define $N:=|\Omega^{h}|$ the size of $\Omega^{h}$, or in other words, the number of points in $\Omega^{h}$.
@@ 4.2 Numerical realization of the total variation
$\begin{split}\min_{u_{\theta}\in\mathcal{H}_{c}^{M}}\Big{\{}E_{\theta}(u_{\theta}):=\alpha_{1}\sum_{x\in\Omega^{h}}w_{1}^{x}|(T^{h}u_{\theta}(x)-g(x)|&+\alpha_{2}\sum_{x\in\Omega^{h}}w_{2}^{x}|(T^{h}u_{\theta}(x)-g(x)|^{2}\\ &+\lambda\sum_{x\in\Omega^{h}}w_{\operatorname{TV}}^{x}|\nabla^{h}u_{\theta}(x)|_{\blacktriangleright}\Big{\}},\end{split}$ (19)
@@ 4.3 $\Gamma$-convergence
$\mathcal{E}_{n}(u):=\begin{cases}\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}+\lambda\int_{\Omega}|\nabla u(x)|\text{d}x&\text{if }u\in\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}^{M_{n}},\\ +\infty&\text{otherwise},\end{cases}$ (24)
where $(c_{n})_{n}\subset\mathbb{R}$ and $(M_{n})_{n}\subset\mathbb{N}$. Recall that for every $c_{n}\geq 0$ and $M_{n}$ the space $\mathcal{H}_{c_{n}}^{M_{n}}$ is a set of continuous piecewise affine functions. If $c_{n}$ and $M_{n}$ are large enough (possibly infinite) then $\mathcal{A}_{h_{n}}\subseteq\mathcal{H}_{c_{n}}^{M_{n}}$ by [7, Theorem 2.1]. Moreover, $0\in\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}^{M_{n}}$ and hence $\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}^{M_{n}}\neq\emptyset$ for any $n\in\mathbb{N}$.
We say that the discrete functional in (19) $\Gamma$-converges to $\mathcal{E}$ defined in (13), if $(\mathcal{E}_{n})_{n}$, where $\mathcal{E}_{n}$ is defined in (24), $\Gamma$-converges to $\mathcal{E}$, cf. [13, Definition 4.2].
@@
Let $\mathcal{E}_{n},\mathcal{E}:L^{1}(\Omega)\to\overline{\mathbb{R}}$ be defined as in (24) and (13), respectively, and $u_{n}$ be a minimizer of $\mathcal{E}_{n}$, $n\in\mathbb{N}$. Then there exist sequences $(c_{n})_{n}\subset\mathbb{R}$ and $(M_{n})_{n}\subset\mathbb{N}$ such that $(\mathcal{E}_{n})_{n}$ $\Gamma$-converges to $\mathcal{E}$ with respect to the $L^{1}(\Omega)$ topology. Moreover, if $\mathcal{E}$ is coercive on $BV(\Omega)$, then we have that
@@
Since $\mathcal{E}$ is coercive on $BV(\Omega)$ it possesses a minimizer in $BV(\Omega)$ [35, Theorem 3.3] and the set $S:=\{u\in BV(\Omega)\colon\mathcal{E}(u)\leq a\}$ is bounded in $BV(\Omega)$, where $a:=\mathcal{E}(0)$. Since $0\in BV(\Omega)$ and $0\in\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}^{M_{n}}$ for all $n\in\mathbb{N}$, any minimizer of $(\mathcal{E}_{n})_{n}$ and $\mathcal{E}$ is in $S$. Thereby theorem 3.4 ensures a minimizer of $\mathcal{E}_{n}$. Since $\overline{S}$ is compact in $L^{1}(\Omega)$ [2, Theorem 2.5] and such that $\inf_{L^{1}(\Omega)}\mathcal{E}_{n}=\inf_{\overline{S}}\mathcal{E}_{n}$, the $\Gamma$-convergence of $(\mathcal{E}_{n})_{n}$ to $\mathcal{E}$ and [13, Theorem 1.21] yield
@@
Contrary to the proof of theorem 3.7 and [34], in the proof of theorem 4.1 we do not use the nested property of the mesh. However, this can be easily achieved by requesting that $c_{n}\geq c_{n-1}$ and $M_{n}\geq M_{n-1}$, whereby at most neurons to the existing hidden layers are added, for all $n>1$, as in the proof of theorem 3.7.
@@
Note that if we set $c_{n}$ and $M_{n}$ in (24) such that $\mathcal{A}_{h_{n}}\cap\mathcal{H}_{c_{n}}^{M_{n}}=\mathcal{A}_{h_{n}}$, then theorem 4.1 shows the $\Gamma$-convergence of a finite difference approximation of $\mathcal{E}$ to $\mathcal{E}$, as in [34]. We emphasize that showing the $\Gamma$-convergence of a finite difference approximation of $\mathcal{E}$ to $\mathcal{E}$ also seems to require a continuous approximation of a $C^{\infty}$-function. In fact, for piecewise constant approximations a $\Gamma$-convergence result cannot be shown for total variation minimization, as a counterexample presented in [10] illustrates.